# Understanding Discrepancies of Wavefunction Theories for Large Molecules

### v1

@@
Abstract
@@
keywords:
Coupled-cluster theory, Many-electron perturbation theory, Theoretical quantum chemistry, First principles, Quantum Monte Carlo1 Introduction
@@
The prediction of electronic transition energies for a single hydrogen atom in 1926 marks the beginning of an incredible successful era of quantum mechanics [1, 2]. Shortly after this breakthrough, Paul Dirac famously noted that the underlying physical laws necessary for much of physics and all of chemistry are now completely known. However, he also pointed out that the exact application of these laws leads to equations that are much too complicated to solve [3]. This paradigm of theoretical chemistry and physics is prevalent until today. In particular, the exponential growth of the computational complexity of the many-electron problem with system size still makes an exact solution of the electronic Schrödinger equation for more than a few atoms impossible. As a consequence, a hierarchy of increasingly accurate methods that is capable of producing reference results at the expense of tractable yet high computational cost has emerged. These reference results are pivotal in order to develop, assess, and further improve computationally more efficient but in general less accurate approximations. In this context, a prime example was the numerical prediction of highly accurate ground state energies for the uniform electron gas using diffusion Monte Carlo methods, leveraging the development of approximate exchange-correlation functionals that ultimately led to the breakthrough of density functional theory in computational materials science during the last decades [4, 5, 6].
@@
At present, quantum mechanical many-electron calculations of systems containing more than 100 atoms have become possible thanks to methodological developments and considerable growth in computing power. These methodological improvements are often based on taking advantage of the relative short-rangedness of many-electron correlation effects [7, 8, 9]. In this manner, the scaling of the computational complexity with respect to system size can be lowered. However, recently several works [10, 11, 12] showed that there exist alarming discrepancies between predicted interaction energies for large molecules when using two of the most widely-trusted highly accurate many-electron theories: DMC and CCSD(T), which stand for diffusion Monte Carlo and coupled-cluster theory using single, double, and perturbative triple particle-hole excitation operators, respectively. These observations are a source of great concern in the electronic structure theory community because, in the case of noncovalent interactions between molecular complexes, both CCSD(T) and DMC are considered highly reliable benchmark methods [13, 14]. Furthermore, the observed discrepancies are large enough to cause qualitative differences in calculated properties of materials, which can have scientific, technological, and even clinical implications. For example, accurate crystal structure predictions are crucial in drug design to differentiate between harmful and effective polymorphs [15, 16, 17]. Similarly, reliable reference methods are essential for discovering and designing new functional materials for applications such as renewable energy storage and conversion, including catalysis, or solar cells [18, 19, 20]. As machine learning increasingly pervades all areas of computational first-principles physics, the accuracy of these reference methods, which provide the training data, becomes even more critical [21, 22, 23].
@@
In the following, we analyze a set of large molecular systems where large discrepancies between approximated versions of DMC and CCSD(T) were observed [10, 11, 12]. Importantly, a direct experimental measurement of the computed interaction energies of these systems is complicated and prone to significant uncertainties. Therefore, it is an open challenge to identify the origin of the observed deviations for the employed highly accurate yet approximate theoretical approaches. A question of paramount importance is, are errors introduced by the fixed-node approximation in DMC and the truncated cluster ansatz of CCSD(T) causing these discrepancies? Or, are approximations used to reduce the scaling of the computational complexity not accurate enough to make reliable predictions?
@@ 2 Results
We present an approach to CCSD(T) calculations of large molecules which builds upon our previous work. This approach enables us to unambiguously test if the employed approximations for DMC and CCSD(T) cause the puzzling discrepancies between their predictions. In particular, our methodology exhibits three striking advantages. Firstly, due to its efficient and massive computational parallelization, we omit any local correlation approximation, as was employed for the CCSD(T) calculations in Refs. [10, 11]. Secondly, we use a plane wave basis set to enable an unbiased assessment of the quality of previously employed tabulated atom-centered basis functions. Thirdly, we are able to study the influence of higher-order contributions to the many-electron perturbation expansion beyond CCSD(T) theory for large molecular complexes.
2.1 Assessing the representation of the wavefunction
In practice, every calculation employing a wavefunction ansatz expands the true wavefunction in a finite basis. Tabulated atom-centered localized Gaussians are the most widely-used basis set for the expansion of many-electron wavefunctions in quantum chemical calculations of large molecules. They achieve a transferable and well-balanced description of electronic correlation effects for a wide range of systems. However, despite their computational advantages, they suffer from one major drawback. As they approach completeness by including more and more diffuse functions, needed to capture polarization effects, their near-linear dependence increases, especially for densely packed structures. Although there exist techniques to ameliorate this problem [24, 25, 26], it remains difficult to employ large basis sets for large systems as done in this work [25].
In contrast to localized basis sets, plane waves form a systematically improvable orthonormal basis and are typically employed in combination with finite simulation cells and periodic boundary conditions. Therefore, computed interaction energies need to be converged with respect to both cell size and plane wave basis set size. Furthermore, compared to the number of Gaussians, the required number of plane wave basis functions is significantly larger, making the latter approach computationally more expensive. To remedy the computational bottlenecks associated with plane wave basis calculations of molecules, we have developed efficient techniques that transform the large plane wave basis set to a system-specific and significantly more compact as well as systematically improvable natural orbital basis [27, 28]. Although the computational cost to determine these transformations is significant, it remains small compared to the computational cost of CCSD(T) calculations for large systems. A detailed describtion of our plane wave approach can be found in the supplementary information.
We now seek to demonstrate that our plane wave basis approach works reliable for the study of noncovalent interactions between molecules and combines the best of two worlds: compactness and systematic improvability without linear dependencies. To this end we turn to the discussion of the computed interaction energy of the parallel displaced benzene dimer on the level of CCSD(T) theory and compare to results from basis set converged Gaussian calculations (see supplementary information).
The total CCSD(T) energy is composed of three terms, the HF total energy, the CCSD correlation energy and the perturbative triples contribution which we denote as (T). Fig. 1(a) depicts the convergence of the CCSD and (T) correlation energy contributions to the computed CCSD(T) interaction energy for a fixed box size with respect to the number of basis functions (natural orbitals) per occupied state (). We include a recently introduced correction to accelerate the convergence of correlation energies to the complete basis set limit (CBS) [29]. Our findings show that a basis set size of suffices to achieve convergence to within a fraction of . We employ this basis set to compute the CCSD(T) interaction energies and its Hartree–Fock (HF), CCSD and (T) correlation energy contributions for different simulation cell sizes. Fig. 1(b) shows that these contributions converge rapidly. Our fully converged estimate of the CCSD(T) interaction energy for the parallel displaced benzene dimer is , which is in excellent agreement with results obtained using Gaussian basis sets of .
@@ 2.1 Assessing the representation of the wavefunction
| LNO-CCSD(T) | -20.6 0.6 | Ref. [10] |
| DLPNO-CCSD(T0) | -20.9 0.4 | Ref. [12] |
| DMC | -18.1(8) | Ref. [10] |
| DMC | -17.5(14) | Ref. [11] |
@@
Having established a reliable and systematically improvable framework to achieve CBS limit estimates of CCSD(T) interaction energies, we now turn to the discussion of results for the parallel displaced coronene dimer (C2C2PD) also studied in Refs. [10, 11]. In this case, significant discrepancies between CCSD(T) and DMC of about after subtracting error bars were reported. The convergence with respect to box size and number of natural orbitals is presented in detail in the supplementary information. Table 1 summarizes the computed interaction energies on different levels of theory. We find that our canonical CCSD(T) estimate agrees to within with domain based local pair-natural orbital (DLPNO-CCSD(T0)) and local natural orbital (LNO-CCSD(T)) findings from Refs. [10, 12]. Therefore, basis set incompleteness errors cannot be the cause for the discrepancy between DMC and CCSD(T) in this case. Likewise, local approximations used for LNO and DLPNO approaches can also be ruled out. However, it is noteworthy that the CCSD(T) interaction energy contains a large (T) contribution of about , indicating that the correct treatment of triple particle-hole excitation effects for the electronic correlation plays a crucial role.
2.2 All that glitters is not gold: overcorrelation of CCSD(T)
The most important question of the present work is: what is the origin of the discrepancy between CCSD(T) and DMC molecular interaction energies? Having ruled out errors from local approximations and incomplete basis sets for the parallel displaced coronene dimer (C2C2PD), we now seek to assess the (T) approximation, which makes a significant contribution to the interaction energy of C2C2PD. In passing we anticipate that the (T) contribution to the interaction energy is also relatively large for all other systems with a significant discrepancy reported in Ref. [10] (see supplementary information).
The (T) approximation was introduced in the seminal work by Raghavachari et. al. [30]. Since then, it has become one of the most widely-used benchmark methods –sometimes referred to as the ’gold standard’ of molecular quantum chemistry– for systems that can be well approximated using single reference based approaches. However, we argue that the partly significantly too strong interaction energies in CCSD(T) theory are caused by the employed truncation of the perturbation series in the approximation of the triple particle-hole excitation operator. As already demonstrated by Nguyen et. al., even weak intermolecular interactions cannot be approximated in a systematically improvable manner by truncated high-order perturbation theories for systems with large polarizabilities [31]. This is also reflected in the too strong interaction energies on the level of MP2 theory, as can be observed for C2C2PD in Table 1. In the extreme case of an infinite polarizability, as it occurs in metallic systems, MP2 and CCSD(T) even yield divergent correlation energies in the thermodynamic limit, which is referred to as infrared catastrophe [32, 33]. In contrast, perturbation theories based on the resummation of certain terms to infinite order can yield interaction energies with an accuracy that is less dependent on the polarizability. Prominent examples for such approaches include the CCSD theory as well as the random-phase approximation. We have recently presented a method, denoted as CCSD(cT), that averts the infrared catastrophe of CCSD(T) by including selected higher-order terms in the triples amplitude approximation without significantly increasing the computational complexity [33].
@@ 2.2 All that glitters is not gold: overcorrelation of CCSD(T)
For the present work it is important to note that the main difference between CCSD(cT) and CCSD(T) theory originates from the employed approximation to the triple particle-hole excitation amplitudes. The triple amplitudes of the (cT) approximation are given in diagrammatic and algebraic form by [33]
@@
where and stand for the Coulomb interaction and the double particle-hole excitation operator, respectively. For brevity, the contributions from the single excitation operator are not included and only one additional ‘direct’ diagram is depicted. In here, , with ’s being one-electron HF energies. The bra- and ket-states correspond to a triple excited and reference state, respectively. The (T) approximation disregards the term , which is included in (cT) and also occurs in full CCSDT theory. This term effectively screens the bare Coulomb interaction of the term and has an opposite sign, making it crucial for systems with large polarizability. As discussed by Nguyen et. al. [31], truncated finite-order perturbation theories can exhibit large errors in noncovalent interactions due to incomplete screening. However, for small and weakly polarizible systems the contribution is small, making the (T) and (cT) approximation agree, as it was already shown for a small set of molecules [33].
@@
We now demonstrate that using CCSD(cT) instead of CCSD(T) theory restores excellent agreement for noncovalent interaction energies with DMC findings. First, we consider again the coronene dimer. Table 1 shows that the binding energy for the coronene dimer calculated on the level of CCSD(cT) theory is by almost closer to the DMC estimate compared to CCSD(T) theory, achieving chemical accuracy () in comparison to DMC after subtracting error bars. Next, we investigate the accuracy of CCSD(T) and CCSD(cT) compared to DMC for noncovalent interactions in smaller molecules. To this end, we study a set of dimers containing up to 24 atoms that were also investigated in Ref. [10]. This gives us another opportunity to assess the effect of local approximations at the level of CCSD(T) theory.
@@
Fig. 2 depicts the deviations of all computed interaction energies from DMC reference values taken from Ref. [10]. It should be noted that DMC references and differences to LNO-CCSD(T) interaction energies are shown with error bars derived from statistical noise and the estimated local approximation error, respectively [10]. Using our massive computational parallelization approach, we are able to add canonical CCSD(T) interaction energies extrapolated to the CBS limit to the comparison to DMC. We stress that for these relatively small molecules, the employed atom-centered Gaussian basis sets together with a standard extrapolation procedure achieve excellent agreement with CBS limit estimates as verified on the level of MP2 theory and summarized in the supplementary information. Importantly, our canonical CCSD(T) results are in good agreement with LNO-CCSD(T) findings to within its error bars. The only minor exception is observed for the parallel displaced uracil dimer, where canonical CCSD(T) predicts a slightly stronger interaction. A comparison to DMC reveals that CCSD(T) theory predicts on average about stronger interaction energies. Based on LNO-CCSD(T) and DMC data alone such a statement cannot be made due to the relatively large and mostly overlapping error bars. However, our well converged canonical CCSD(T) findings allow drawing such conclusions. Only for the T-shaped pyridine and benzene dimers, DMC and CCSD(T) binding energies agree to within the DMC errors. Note that these systems have a smaller (T) contribution to the intereaction energy, compared to the parallel displaced systems. All other systems exhibit small but significant discrepancies between CCSD(T) and DMC results, which is consistent with the even larger discrepancies reported for the larger molecules in Ref. [10]. Similar to our findings for the coronene dimer reported in Table 1, Fig. 2 shows that CCSD(cT) interaction energies agree significantly better with DMC values than their CCSD(T) counterparts.
Given the good agreement between DMC and CCSD(cT) for the systems studied above, an important question to ask is if CCSD(cT) is really more accurate than CCSD(T) for noncovalent interaction energies? To answer this question, following the paradigm of theoretical quantum chemistry, we turn to more accurate reference methods. CC methods form a rapidly convergent series of approximations to the true ground state wavefunction and energy, especially for weakly interacting systems [37]. This implies that the following hierarchy applies to absolute energy differences: CCSD-HF CCSDT-CCSD CCSDTQ-CCSDT. Therefore, we first seek to assess if CCSD(cT) is a better approximation to CCSDT than CCSD(T) theory. Our massive computational parallelization approach allows us to compute CCSDT correlation energies for a set of molecules contained in the S22 set employing cc-pVDZ basis sets [34], which suffices for the following qualitative discussion. The respective CCSDT and CCSD(T) correlation energies are depicted in Fig. 3(a). Our findings show that CCSD(T) and CCSDT energies are in good agreement even in the case of large molecular complexes. However, in contrast to total correlation energies, a comparison between CCSDT and CCSD(T) interaction energies of the same molecular complexes draws a completely different picture. Fig. 3(b) shows that the larger the triples contribution to the intermolecular interaction energy gets in magnitude, the larger CCSD(T) overestimates compared to CCSDT. We suggest the following conclusion: CCSD(T) and CCSDT total correlation energies agree fortuitously well, which fails in the case of physically more relevant strong interaction energies. The opposite applies to CCSD(cT), which agrees less good with CCSDT for absolute correlation energies, but very well in the case of interaction energies depicted in Figs. 3(a) & 3(b), respectively.

### v3

@@
Understanding Discrepancies of Wavefunction Theories for Large Molecules
@@
1 Introduction
@@
The prediction of electronic transition energies for a single hydrogen atom in 1926 marks the beginning of an incredible successful era of quantum mechanics (?, ?). Shortly after this breakthrough, Paul Dirac famously noted that the underlying physical laws necessary for much of physics and all of chemistry are now completely known. However, he also pointed out that the exact application of these laws leads to equations that are much too complicated to solve (?). This paradigm of theoretical chemistry and physics is prevalent until today. In particular, the exponential growth of the computational complexity of the many-electron problem with system size still makes an exact solution of the electronic Schrödinger equation for more than a few atoms impossible. As a consequence, a hierarchy of increasingly accurate methods that is capable of producing reference results at the expense of tractable yet high computational cost has emerged. These reference results are pivotal in order to develop, assess, and further improve computationally more efficient but in general less accurate approximations. In this context, a prime example was the numerical prediction of highly accurate ground state energies for the uniform electron gas using diffusion Monte Carlo methods, leveraging the development of approximate exchange-correlation functionals that ultimately led to the breakthrough of density functional theory in computational materials science during the last decades (?, ?, ?).
@@
At present, quantum mechanical many-electron calculations of systems containing more than 100 atoms have become possible thanks to methodological developments and considerable growth in computing power. These methodological improvements are often based on taking advantage of the relative short-rangedness of many-electron correlation effects (?, ?, ?). In this manner, the scaling of the computational complexity with respect to system size can be lowered. However, recently several works (?, ?, ?, ?) showed that there exist alarming discrepancies between predicted interaction energies for large molecules when using two of the most widely-trusted highly accurate many-electron theories: DMC and CCSD(T), which stand for diffusion Monte Carlo and coupled-cluster theory using single, double, and perturbative triple particle-hole excitation operators, respectively. These observations are a source of great concern in the electronic structure theory community because, in the case of noncovalent interactions between molecular complexes, both CCSD(T) and DMC are considered highly reliable benchmark methods (?, ?). Furthermore, the observed discrepancies are large enough to cause qualitative differences in calculated properties of materials, which can have scientific, technological, and even clinical implications. For example, accurate crystal structure predictions are crucial in drug design to differentiate between harmful and effective polymorphs (?, ?, ?). Similarly, reliable reference methods are essential for discovering and designing new functional materials for applications such as renewable energy storage and conversion, including catalysis, or solar cells (?, ?, ?). Finally, as machine learning increasingly pervades all areas of computational first-principles physics, the accuracy of these reference methods, which provide the training data, becomes even more critical (?, ?, ?).
@@
In the following, we analyze a set of large molecular systems where large discrepancies between approximated versions of DMC and CCSD(T) were observed (?, ?). Importantly, a direct experimental measurement of the computed interaction energies of these systems is complicated and prone to significant uncertainties. Therefore, it is an open challenge to identify the origin of the observed deviations for the employed highly accurate yet approximate theoretical approaches.
@@ 2 Results
We present an approach which allows us to unambiguously test if the employed approximations for DMC and CCSD(T) cause the puzzling discrepancies between their predictions. In particular, our methodology exhibits three striking advantages. Firstly, due to its efficient and massive computational parallelization, we omit any local correlation approximation, as was employed for the CCSD(T) calculations in Refs. (?, ?, ?). Secondly, we use a plane wave basis set to enable an unbiased assessment of the quality of previously employed tabulated atom-centered basis functions. Thirdly, we are able to study the influence of higher-order contributions to the many-electron perturbation expansion beyond CCSD(T) theory for large molecular complexes.
In order to demonstrate the reliability of our plane wave basis approach, we first investigate the parallel displaced benzene dimer as a benchmark. We find that our approach effectively addresses the challenges of noncovalent interactions between large molecules, combining the compactness and systematic improvability of natural orbitals without near-linear dependencies that plague atom-centered Gaussian basis sets for densely packed structures. As discussed in the supplementary information, our computed CCSD(T) interaction energies for the parallel displaced benzene dimer are in excellent agreement with Gaussian basis set results. Next, we turn to the coronene dimer interaction energy, where significant discrepancies between DMC and CCSD(T) have been observed (?, ?). As shown in Tab. 1, our canonical CCSD(T) estimates for the parallel displaced coronene dimer align closely with domain-based local pair-natural orbital (DLPNO-CCSD(T)) and local natural orbital (LNO-CCSD(T)) results, ruling out basis set incompleteness and local approximation errors as sources of discrepancies with DMC findings. However, it is noteworthy that the CCSD(T) interaction energy contains a large (T) contribution of about , indicating that the correct treatment of triple particle-hole excitation effects for the electronic correlation plays a crucial role.
@@
LNO-CCSD(T) | -20.6 0.6 | Ref. (?) |
DLPNO-CCSD(T0) | -20.9 0.4 | Ref. (?) |
DMC | -18.1(8) | Ref. (?) |
DMC | -17.5(14) | Ref. (?) |
@@
2.1 All that glitters is not gold: overcorrelation in CCSD(T)
Having ruled out errors from local approximations and incomplete basis sets for the parallel displaced coronene dimer (C2C2PD), we seek to assess the (T) approximation, which makes a significant contribution to the interaction energy of C2C2PD. In passing we anticipate that the (T) contribution to the interaction energy is also relatively large for all other systems with a significant discrepancy reported in Ref. (?) (see supplementary information).
The (T) approximation was introduced in the seminal work by Raghavachari et. al. (?). Since then, it has become one of the most widely-used benchmark methods –sometimes referred to as the ’gold standard’ of molecular quantum chemistry– for weakly correlated systems. However, we argue that the partly significantly too strong interaction energies in CCSD(T) theory are caused by the employed truncation of the approximation of the triple particle-hole excitation operator. These shortcomings are comparable to the issue of too strong interaction energies from truncated perturbation theories for systems with large polarizabilities, as discussed by Nguyen et. al. (?). As can be observed for the coronene dimer in Table 1, second-order Møller-Plesset perturbation theory (MP2)—a truncated pertubation theory—exhibits this overestimation of the interaction energy. In the extreme case of an infinite polarizability, as it occurs in metallic systems, MP2 and CCSD(T) even yield divergent correlation energies in the thermodynamic limit, which is referred to as infrared catastrophe (?, ?). In contrast, a resummation of certain terms to infinite order can yield interaction energies with an accuracy that is less dependent on the polarizability. Prominent examples for such approaches include the CCSD theory as well as the random-phase approximation. We have recently presented a method, denoted as CCSD(cT), that averts the infrared catastrophe of CCSD(T) by including selected higher-order terms in the triples amplitude approximation without significantly increasing the computational complexity (?).
@@ 2.1 All that glitters is not gold: overcorrelation in CCSD(T)
For the present work it is important to note that the main difference between CCSD(cT) and CCSD(T) theory originates from the employed approximation to the triple particle-hole excitation amplitudes. The triple amplitudes of the (cT) approximation are given in diagrammatic and algebraic form by (?)
@@
where and stand for the Coulomb interaction and the double particle-hole excitation operator, respectively. For brevity, the contributions from the single excitation operator are not included and only one additional ‘direct’ diagram is depicted. In here, , with ’s being one-electron HF energies. The bra- and ket-states correspond to a triple excited and reference state, respectively. The (T) approximation disregards the term , which is included in (cT) and also occurs in full CCSDT theory. This term effectively screens the bare Coulomb interaction of the term and has an opposite sign, making it crucial for systems with large polarizability. However, for small and weakly polarizible systems the contribution is small, making the (T) and (cT) approximation agree, as it was already shown for a set of small molecules (?).
@@
We now demonstrate that using CCSD(cT) instead of CCSD(T) theory restores excellent agreement for noncovalent interaction energies with DMC findings. First, we consider again the coronene dimer. Table 1 shows that the binding energy for the coronene dimer calculated on the level of CCSD(cT) theory is by almost closer to the DMC estimate compared to CCSD(T) theory, achieving chemical accuracy () in comparison to DMC after subtracting error bars. Next, we investigate the accuracy of CCSD(T) and CCSD(cT) compared to DMC for noncovalent interactions in smaller molecules. To this end, we study a set of dimers containing up to 24 atoms that were also investigated in Ref. (?). This gives us another opportunity to assess the effect of local approximations at the level of CCSD(T) theory.
@@
Fig. 1 depicts the deviations of all computed interaction energies from DMC reference values taken from Ref. (?). It should be noted that DMC references and differences to LNO-CCSD(T) interaction energies are shown with error bars (?). Using our massive computational parallelization approach, we are able to add canonical CCSD(T) interaction energies extrapolated to the CBS limit to the comparison to DMC. For these relatively small molecules, we can employ sufficiently large basis sets, reducing the remaining uncertainty to approximately 0.01 kcal/mol (see supplementary information). Importantly, our canonical CCSD(T) results are in good agreement with LNO-CCSD(T) findings to within its error bars. The only minor exception is observed for the parallel displaced uracil dimer, where canonical CCSD(T) predicts a slightly stronger interaction. A comparison to DMC reveals that CCSD(T) theory predicts on average about stronger interaction energies. Based on LNO-CCSD(T) and DMC data alone such a statement cannot be made due to the relatively large and mostly overlapping error bars. However, our well converged canonical CCSD(T) findings allow drawing such conclusions. Only for the T-shaped pyridine and benzene dimers, DMC and CCSD(T) binding energies agree to within the DMC errors. Note that these systems have a smaller (T) contribution to the intereaction energy, compared to the parallel displaced systems. All other systems exhibit small but significant discrepancies between CCSD(T) and DMC results, which is consistent with the even larger discrepancies reported for the larger molecules in Ref. (?). Similar to our findings for the coronene dimer reported in Table 1, Fig. 1 shows that CCSD(cT) interaction energies agree significantly better with DMC values than their CCSD(T) counterparts.
Given the good agreement between DMC and CCSD(cT) for the systems studied above, an important question to ask is if CCSD(cT) is really more accurate than CCSD(T) for noncovalent interaction energies? To answer this question we compare interaction energies of both approaches to higher-level CC methods for complexes from the S22 data set. As can be observed in Fig. 2, we find that while (T) is in good agreement with T for total energies, it overestimates interaction energies. Here T stands for the triples contribution to the correlation energy, . (cT) closely matches the T interaction energies, indicating its superior accuracy for weakly bound complexes. This effect is particularly strong for interaction energies with large triples correlation contributions.